OpenStudy (anonymous):

Simplify 8xy^3 z divided by the quantity 2^-2 x raised to the -3 power y raised to the zero power z. 32x^3y^3 8x^4y^3 32x^4y^3 32x^4y^2

OpenStudy (anonymous):

Are you looking to simplify \[\frac{8xy^{3}z}{2^{-2}x^{-3}y^0 z} ?\]

OpenStudy (anonymous):

yes that's it :) @KevinOrr

OpenStudy (anonymous):

Do you know what negative exponents do?

OpenStudy (anonymous):

No im not so sure @KevinOrr

OpenStudy (anonymous):

\[x^{-k} = \frac{1}{x^k}\\ \frac{1}{x^{-k}} = x^k\]

OpenStudy (anonymous):

So, since \(2^{-2} = \frac{1}{2^2} = \frac{1}{4}, \frac{1}{x^{-2}} = x^2, \textrm{and} y^0 = 1\), then\[\frac{8xy^3z}{2^{-2}x^{-2}y^0z} = \frac{8 \cdot 4 \cdot x \cdot x^2 \cdot y^3 \cdot z}{1 \cdot z} = 32x^3y^3\]